# Semantic-Enhanced Time-Series Forecasting via Large Language Models

### v3

@@ Methods
Table 1: The input length of the SE-LLM is 672. The forecast horizon is $\{96,192,336,720\}$. The best results are in bold and the second best are underlined.
Short-Term Forecasting. It refers to predicting target variables within a relatively brief future time horizon. We have meticulously selected algorithms suitable for short-term forecasting for comparison. SE-LLM was compared with the SOTA models in short-term forecasting on the M4 dataset, as shown in Table 2, and the results demonstrated that we achieved the best performance. Our method achieved a 0.26% reduction in Symmetric Mean Absolute Percentage Error (SMAPE), a 0.94% decline in Mean Absolute Scaled Error (MASE), and a 0.47% decrease in Quantile-Weighted Absolute Error (QWA) relative to the second best.
Models SE-LLM TimeMixer++ DECA Time-VLM AutoTimes S2IP-LLM Time-LLM FPT iTransformer Dlinear PatchTST TimesNet
2025 2025 2025 2024 2024 2024 2024 2024 2023 2023 2023
Yearly SMAPE 13.294 13.397 13.288 13.419 13.319 13.413 13.419 15.531 13.652 14.012 13.517 13.394
MASE 2.970 2.990 2.974 3.005 3.021 3.024 3.005 3.015 3.095 3.071 0.301 3.004
OWA 0.780 0.786 0.781 0.789 0.789 0.792 0.789 0.793 0.807 0.815 0.795 0.787
Quarterly SMAPE 10.079 10.206 10.037 10.110 10.020 10.352 10.110 10.177 10.353 10.758 10.847 10.101
MASE 1.177 1.201 1.174 1.178 1.162 1.228 1.178 1.194 1.209 1.306 1.315 1.183
OWA 0.887 0.901 0.884 0.889 0.878 0.922 0.889 0.897 0.911 0.905 0.972 0.890
Monthly SMAPE 12.618 12.720 12.762 12.980 12.696 12.995 12.980 12.894 13.079 13.377 14.854 12.866
MASE 0.931 0.943 0.947 0.963 0.936 0.970 0.963 0.956 0.974 1.021 1.169 0.964
OWA 0.875 0.884 0.897 0.903 0.880 0.910 0.903 0.897 0.911 0.944 1.055 0.894
Others SMAPE 4.896 4.593 4.761 4.795 4.916 4.805 4.795 4.940 4.780 5.259 6.184 4.892
MASE 3.306 3.380 3.207 3.178 3.310 3.247 3.178 3.228 3.231 3.608 4.818 3.323
OWA 1.037 1.054 1.007 1.006 1.039 1.017 1.006 1.029 1.012 1.122 1.140 1.048
Average SMAPE 11.778 11.884 11.828 11.894 11.831 12.021 11.983 11.991 12.142 12.489 13.022 11.930
MASE 1.578 1.597 1.580 1.592 1.585 1.612 1.595 1.600 1.631 1.690 1.814 1.597
OWA 0.847 0.856 0.850 0.855 0.850 0.857 0.859 0.861 0.874 0.902 0.954 0.867
Zero-Shot forecasting. The TSCC module employs a VAE to construct simulated noisy data, enabling the learning of latent distribution characteristics in time series and generating diverse synthetic samples. This approach enhances the model’s generalization capability for unseen data in zero-shot prediction scenarios, making it suitable for forecasting previously unencountered temporal patterns. We conducted zero-shot prediction experiments on the univariate datasets M3 and M4. Since most existing studies are not suitable for generalization on M3 and M4, we selected the SOTA methods listed in Table 3 for comparison. Specifically, the M3 and M4 datasets exhibit various temporal variation patterns with different data distributions. We discuss the domain generalization between these datasets. The results indicate that SE-LLM achieved the best performance in the zero-shot generalization comparison, further demonstrating its generalization capability. Compared to AutoTimes, the SMAPE decreased by 0.1% when transferring from M3 to M4, and by 1.4% when transferring from M4 to M3.
Others 6.269 6.259 7.081 9.194 7.087 6.863 7.529 7.348 7.001
Avg 13.024 13.036 13.125 15.337 13.228 14.553 15.047 19.047 14.092
Quarterly 9.011 9.350 10.130 9.740 9.620 12.260 9.480 13.000 9.760
Others 4.872 5.790 4.810 6.810 9.440 6.880 8.940 13.030 7.530
Avg 12.560 12.750 13.060 14.030 13.390 14.170 13.530 15.820 13.370
In-Context Forecasting. It was first introduced in AutoTimes (Liu et al. 2024b). Similar to AutoTime, we employ an autoregressive forecasting approach, enabling In-Context Forecasting. To rigorously evaluate SE-LLM’s performance, we conducted a comparative study against this baseline. The proposed approach shares conceptual similarities with zero-shot prediction, wherein a model trained on source domain data $\mathcal{D}_{s}$ is directly applied to target domains $\mathcal{D}_{t}$ without retraining. The experimental results are shown in Fig. 5, indicating that we have achieved better performance.
Ablation Study
Ablation on Innovation Methods. Ablation experiments are conducted on the ECL and Traffic datasets. To reduce computational costs, we selected different lightweight LLMs for comparison, including GPT2 (Radford et al. 2019), BERT (Devlin et al. 2019), Opt-125m (Zhang et al. 2022), and Qwen2.5-0.5B (Yang et al. 2024a), which are commonly used in current research (Zhou et al. 2023; Liu et al. 2024b, 2025; Jin et al. 2024). The experimental results are shown in Table 4. Our baseline uses a cross-domain attention mechanism for multi-modal alignment (Chang, Peng, and Chen 2025; Jin et al. 2024; Liu et al. 2025). After replacing it with the TSCC framework and adding the Time-Adapter, the prediction accuracy of different LLMs has improved, demonstrating the effectiveness of the innovative methods in this paper. Comparative experiments show Qwen-0.5B yields lower prediction errors among LLMs.
Figure 5: Results on M4 dataset. ”Others” denotes the average values for weeks, days, and hours.
LLM GPT2 Bert Opt125m Qwen2.5-0.5B
Dataset ELC Traffic ELC Traffic ELC Traffic ELC Traffic
Mertic MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Baseline 0.145 0.240 0.387 0.267 0.150 0.245 0.440 0.304 0.140 0.237 0.368 0.258 0.135 0.233 0.373 0.263
0.165 0.257 0.410 0.277 0.168 0.262 0.459 0.314 0.158 0.254 0.390 0.268 0.152 0.249 0.394 0.272
0.183 0.275 0.427 0.286 0.187 0.281 0.476 0.322 0.176 0.272 0.407 0.292 0.170 0.267 0.409 0.280
0.224 0.309 0.463 0.306 0.230 0.316 0.514 0.344 0.216 0.305 0.445 0.300 0.212 0.303 0.444 0.300
0.179 0.270 0.422 0.284 0.184 0.276 0.472 0.321 0.173 0.267 0.403 0.280 0.167 0.263 0.405 0.279
+TSCC 0.138 0.233 0.368 0.257 0.144 0.240 0.395 0.277 0.136 0.232 0.361 0.251 0.134 0.231 0.356 0.247
0.157 0.250 0.393 0.269 0.161 0.256 0.417 0.286 0.154 0.249 0.383 0.263 0.152 0.248 0.377 0.257
0.176 0.269 0.414 0.281 0.181 0.274 0.436 0.295 0.173 0.267 0.401 0.271 0.169 0.266 0.393 0.266
0.218 0.305 0.461 0.312 0.223 0.309 0.474 0.316 0.215 0.303 0.439 0.294 0.210 0.301 0.431 0.287
0.172 0.264 0.409 0.280 0.177 0.270 0.431 0.294 0.170 0.263 0.396 0.270 0.166 0.262 0.389 0.264
+Time-Adapter 0.133 0.227 0.363 0.255 0.135 0.230 0.378 0.266 0.132 0.227 0.354 0.247 0.130 0.225 0.350 0.243
0.152 0.244 0.388 0.267 0.152 0.247 0.400 0.275 0.149 0.243 0.378 0.258 0.148 0.242 0.373 0.254
0.169 0.262 0.411 0.281 0.170 0.265 0.417 0.283 0.166 0.261 0.396 0.267 0.164 0.259 0.390 0.263
0.211 0.299 0.461 0.314 0.209 0.298 0.454 0.303 0.204 0.295 0.434 0.291 0.203 0.293 0.429 0.285
0.166 0.258 0.406 0.279 0.167 0.260 0.412 0.282 0.163 0.257 0.391 0.266 0.161 0.255 0.386 0.261
Ablation on Adapters. Long-term forecasting should capture long-term dependencies, making the task more challenging. Therefore, to demonstrate the performance of Time-Adapter, we conducted a comparison with the LoRA (Hu et al. 2022; Liu et al. 2024b), and the results are presented in Table 5. The experimental baseline is the Qwen2.5 with the TSCC module. Experiments findings clearly demonstrate that LoRA lacks generalization capability to enhance forecasting performance. In contrast, Time-Adapter endows LLM with the ability to adapt to forecasting tasks.
Datasets Traffic ECL Solar
Metrics MSE MAE MSE MAE MSE MAE
Baseline: +TSCC 0.389 0.264 0.166 0.262 0.201 0.251
+LoRA 0.390 0.268 0.163 0.257 0.202 0.253
+Time-Adapter 0.386 0.261 0.161 0.255 0.192 0.243
Table 5: All reported results are averaged. (Full results are in Appendix)
Ablation on LLM-based SOTA methods. The modal disparity between time-series and language data is a challenging problem. To address this issue, we integrated the TSCC and the Time-Adapter as a plugin into the Time-LLM (Jin et al. 2024), AutoTimes (Liu et al. 2024b), and TimeCMA (Liu et al. 2025). We analyzed their impact on algorithm performance using the ETTh and ETTm datasets. Our experiments were conducted based on the parameters provided in the source code. All other experimental parameters remained unchanged when incorporating the TSCC and Time-Adapter, with the GPT-2 model consistently used as the LLM. The experimental results, listed in Table 6, indicate that errors were reduced across the board. However, Time-LLM yielded negligible performance improvements on the ETTm dataset.
Notably, for the TimeCMA model, the Mean Squared Error (MSE) reduces significantly. Specifically, on the ETTh2 dataset, the MSE was reduced by 8.9%. TimeCMA uses large language models as the output of time-series text prompts. The LLM does not participate in the training, making conducting ablation experiments on Time-Adapter impossible. The experiment results highlight that our research addresses a common issue in existing studies.
Datasets ETTh1 ETTh2 ETTm1 ETTm2
Mertics MSE MAE MSE MAE MSE MAE MSE MAE
Time-LLM Baseline 0.420 0.439 0.370 0.404 0.358 0.389 0.262 0.324
TSCC 0.407 0.429 0.368 0.401 0.360 0.390 0.267 0.328
Time-Adapter 0.405 0.428 0.364 0.399 0.357 0.388 0.266 0.326
AutoTimes Baseline 0.397 0.425 0.367 0.410 0.361 0.389 0.274 0.329
TSCC 0.390 0.421 0.363 0.406 0.354 0.384 0.269 0.325
Time-Adapter 0.388 0.420 0.357 0.407 0.351 0.382 0.268 0.323
Time-CMA Baseline 0.446 0.446 0.416 0.429 0.399 0.413 0.294 0.337
TSCC 0.437 0.434 0.379 0.403 0.383 0.396 0.285 0.322
Table 6: The forecasting horizons are in $\{96,192,336,720\}$, and all experimental results presented are averages. (Full results are in Appendix)
Qualitative Analysis
In this section, we conduct a qualitative analysis of how TSCC injects temporal patterns into the joint space. Fig. 6 (a) presents the visualization results of the semantic space, while Fig. 6 (b) displays the learned time series embeddings. Fig. 6 (c) shows the extracted results of anomaly parts in the joint space. The TSCC module accurately reveals distinct clustering characteristics in the data exhibiting anomaly patterns. This indicates that through the TSCC module, we can obtain enhanced semantics with temporal patterns that are more distinguishable.
Figure 6: The t-SNE plots of embeddings space
Computational Efficiency
From the perspectives of computational efficiency and prediction error, the results in Fig. 7 indicate that Qwen2.5-0.5B offers relatively lower computational consumption with lower errors on the ECL dataset, making it more suitable for long-term forecasting in SE-LLM.
Figure 7: The size of the circles reflects the proportional relationship among the parameter counts of different LLMs
Compared to current research, we selected two representative studies, AutoTimes and Time-LLM. First, we compared the performance under their respective optimal prediction conditions. Then, we further evaluated their computational efficiency, including training and inference time. Referring to the results in Fig. 8, our method achieves improved performance.
Figure 8: All experiments use a batch size of 256.
Conclusion
Appendix
Autoregressive Forecasting
In time series forecasting, the autoregressive model can accomplish predictions for any time step length by merely training on data with a fixed step length, significantly reducing computational costs.
Autoregressive models generate future predictions by iteratively conditioning on historical observations. This strictly sequential prediction strategy relies on the temporal continuity and internal patterns embedded in past data. When such patterns remain consistent over time, autoregressive approaches typically yield stable and reliable forecasts.
SE-LLM adopts an autoregressive inference scheme, inspired by the design of AutoTimes. Specifically, we generate future values by recursively feeding the model’s previous outputs into the input sequence. The complete procedure is outlined in Algorithm 2.
Algorithm 2 Gated Attention Prediction Procedure
0: Input time series $\mathbf{X}\in\mathbb{R}^{B\times L}$, where $B$ is the batch size and $L$ is the sequence length. Output sequence $\mathbf{Y}\in\mathbb{R}^{B\times T}$.
1:if $N=T\,\text{div}\,\tau$ then
2:  for $\text{id}\leftarrow 1$ to NN do
3:   $\hat{\mathbf{Y}}\leftarrow\text{Model}(\mathbf{X})$
4:   $\hat{\mathbf{X}}\leftarrow\text{Cat}(\mathbf{X}_{\tau:L},\hat{\mathbf{Y}}_{L-\tau:L})$
5:  end for
6:  return $\hat{\mathbf{X}}$
7:end if
[…]
9:for $\text{id}\leftarrow 1$ to $(N+1)$ do
10:  $\hat{\mathbf{Y}}\leftarrow\text{Model}(\mathbf{X})$
11:  $\hat{\mathbf{X}}\leftarrow\text{Cat}(\mathbf{X}_{\tau:L},\hat{\mathbf{Y}}_{L-\tau:L})$
12:  if $\text{id}=N+1$ then
13:   $\hat{\mathbf{X}}\leftarrow\text{Cat}(\mathbf{X}_{\tau:L},\hat{\mathbf{Y}}_{L-\tau+l:L})$
14:   return $\hat{\mathbf{X}}$
15:  end if
16:end for
17:return $\hat{\mathbf{X}}$
Evalution Metrics
In time series forecasting, quantitative evaluation is essential for assessing model performance. In this study, we adopt standard metrics tailored to different forecasting scenarios. For long-term forecasting, we use Mean Squared Error (MSE) and Mean Absolute Error (MAE):
@@ Conclusion
Time series datasets consist of sequential data points collected or recorded at consistent time intervals (e.g., hourly, daily, monthly). Each entry includes a timestamp and one or more observed values. These datasets are used to train models that identify historical patterns to forecast future values. Common applications include predicting stock prices, energy demand, weather, sales, and website traffic. The detailed content are shown in Table 7.
Tasks Datasets Dim Series Length Dataset Size Frequency Information
Weather 21 $\{96,192,336,720\}$ $\{36792,5271,10540\}$ 10min Weather
ECL 321 $\{96,192,336,720\}$ $\{18317,2633,5261\}$ Hourly Eelectricity
Solar 137 $\{96,192,336,720\}$ $\{36601,5161,10417\}$ 10min Energy
Short-Term Forecasting and Zero-Shot Forecasting M4-Yearly 1 6 $\{23000,0,23000\}$ Yearly Demographic
M4-Quarterly 1 8 $\{24000,0,24000\}$ Quarterly Finace
M4-Monthly 1 18 $\{48000,0,48000\}$ Monthly Industry
M4-Weekly 1 13 $\{359,0,359\}$ Weekly Macro
M4-Daily 1 14 $\{4227,0,4227\}$ Daily Macro
M4-Hourly 1 48 $\{414,0,414\}$ Hourly Other
Zero-Shot Forecasting M3-Yearly 1 6 $\{645,0,645\}$ Yearly Demographic
M3-Quarterly 1 8 $\{756,0,756\}$ Quarterly Finace
M3-Monthly 1 18 $\{1428,0,1428\}$ Monthly Industry
M3-Others 1 8 $\{174,0,174\}$ Others Macro
Full Experiment Resullts
Ablation on Adapters. Full results are shown in Table 8.
@@
Mertics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
TSCC 0.352 0.393 0.162 0.245 0.356 0.247 0.134 0.231 0.175 0.230
0.378 0.411 0.209 0.259 0.377 0.257 0.152 0.248 0.194 0.245
0.391 0.420 0.263 0.299 0.393 0.266 0.169 0.266 0.209 0.257
0.402 0.437 0.332 0.346 0.431 0.287 0.210 0.301 0.227 0.270
0.381 0.415 0.242 0.287 0.389 0.264 0.166 0.262 0.201 0.251
+LoRA 0.362 0.399 0.157 0.210 0.357 0.251 0.132 0.228 0.173 0.228
0.393 0.420 0.204 0.254 0.378 0.260 0.149 0.244 0.194 0.245
0.410 0.432 0.260 0.297 0.394 0.269 0.166 0.261 0.210 0.260
0.418 0.447 0.339 0.351 0.430 0.290 0.205 0.295 0.229 0.277
0.396 0.425 0.240 0.278 0.390 0.268 0.163 0.257 0.202 0.253
+Time-Adapter 0.355 0.398 0.150 0.200 0.350 0.243 0.130 0.225 0.165 0.220
0.389 0.421 0.194 0.243 0.373 0.254 0.148 0.242 0.183 0.236
0.409 0.435 0.247 0.285 0.390 0.263 0.164 0.259 0.199 0.249
0.418 0.448 0.323 0.339 0.429 0.285 0.203 0.293 0.219 0.266
0.393 0.426 0.229 0.267 0.386 0.261 0.161 0.255 0.192 0.243
Table 8: Time-Adapter compared with LoRA. The forecasting horizons were set to $\{96,192,336,720\}$.
Ablation on LLM-based SOTA methods. Full results are shown in Table 9.
Time-LLM
ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
Baseline 0.384 0.411 0.293 0.349 0.300 0.356 0.172 0.265
0.414 0.430 0.379 0.400 0.340 0.378 0.232 0.306
0.433 0.448 0.381 0.414 0.370 0.395 0.281 0.339
0.448 0.468 0.428 0.452 0.422 0.428 0.362 0.387
0.420 0.439 0.370 0.404 0.358 0.389 0.262 0.324
TSCC 0.373 0.401 0.287 0.352 0.302 0.354 0.180 0.270
0.403 0.425 0.367 0.393 0.341 0.378 0.247 0.314
0.414 0.430 0.390 0.414 0.373 0.399 0.276 0.335
0.436 0.460 0.426 0.446 0.422 0.427 0.366 0.391
0.407 0.429 0.368 0.401 0.360 0.390 0.267 0.328
Time-Adapter 0.368 0.400 0.294 0.347 0.303 0.354 0.178 0.268
0.401 0.421 0.370 0.400 0.345 0.376 0.229 0.303
0.412 0.430 0.373 0.402 0.366 0.394 0.291 0.342
0.439 0.459 0.418 0.445 0.414 0.428 0.364 0.391
0.405 0.428 0.364 0.399 0.357 0.388 0.266 0.326
AutoTimes
ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
Baseline 0.360 0.397 0.288 0.352 0.291 0.347 0.176 0.264
0.391 0.419 0.351 0.395 0.338 0.376 0.238 0.307
0.408 0.432 0.383 0.424 0.376 0.398 0.298 0.346
0.429 0.452 0.447 0.468 0.438 0.433 0.382 0.400
0.397 0.425 0.367 0.410 0.361 0.389 0.274 0.329
TSCC 0.354 0.394 0.288 0.352 0.286 0.341 0.174 0.261
0.385 0.415 0.350 0.393 0.334 0.371 0.236 0.303
0.401 0.427 0.378 0.419 0.370 0.394 0.291 0.341
0.420 0.446 0.435 0.461 0.427 0.429 0.373 0.394
0.390 0.421 0.363 0.406 0.354 0.384 0.269 0.325
Time-Adapter 0.352 0.393 0.285 0.377 0.285 0.340 0.173 0.262
0.383 0.414 0.351 0.390 0.332 0.369 0.234 0.302
0.399 0.425 0.376 0.413 0.366 0.392 0.291 0.337
0.417 0.446 0.416 0.446 0.421 0.426 0.373 0.389
0.388 0.420 0.357 0.407 0.351 0.382 0.268 0.323
Time-CMA
ETTh1 ETTh2 ETTm1 ETTm2
MSE MAE MSE MAE MSE MAE MSE MAE
Baseline 0.393 0.414 0.339 0.378 0.340 0.380 0.191 0.274
0.435 0.437 0.421 0.425 0.379 0.407 0.257 0.316
0.472 0.453 0.452 0.452 0.405 0.413 0.313 0.350
0.482 0.479 0.452 0.462 0.472 0.450 0.416 0.407
0.446 0.446 0.416 0.429 0.399 0.413 0.294 0.337
TSCC 0.377 0.396 0.299 0.347 0.311 0.352 0.173 0.256
0.429 0.424 0.377 0.395 0.360 0.381 0.242 0.301
0.468 0.444 0.415 0.426 0.396 0.405 0.307 0.323
0.474 0.471 0.424 0.443 0.464 0.444 0.416 0.407
0.437 0.434 0.379 0.403 0.383 0.396 0.285 0.322

### v4

@@ Appendix A Appendix of SE-LLM
Appendix B Experiment
Experimental Setup
Evaluation Metrics. Time series forecasting, quantitative evaluation is essential for assessing model performance. In this study, we adopt standard metrics tailored to different forecasting scenarios. For long-term forecasting, we use Mean Squared Error (MSE) and Mean Absolute Error (MAE):
@@ Appendix B Experiment
Datasets. We adopted publicly available benchmarks in four research domains: long-term forecasting, short-term forecasting, zero-shot forecasting and in-context forecasting (Liu et al. 2024c). The experimental datasets comprised: ETTh (Zhou et al. 2021), ETTm (Zhou et al. 2021), Weather (Chen et al. 2021), Elcetricity (ECL) (Chen et al. 2021), Traffic (Chen et al. 2021), Solar (Lai et al. 2018), M3 (Wu et al. 2023), and M4 (Makridakis, Spiliotis, and Assimakopoulos 2020) datasets. The detailed introduction refers to the Table 8. Time series datasets consist of sequential data points collected or recorded at consistent time intervals (e.g., hourly, daily, monthly). Each entry includes a timestamp and one or more observed values. These datasets are used to train models that identify historical patterns to forecast future values. Common applications include predicting stock prices, energy demand, weather, sales, and website traffic.
Baseline. The experiments were conducted and compared with SOTA methods, such as: TimeMixer++ (Wang et al. 2025) and AutoTimes (Liu et al. 2024c) as the high-performance baselines for evaluation. The algorithms involved in all the experiments are TimeMOE (Shi et al. 2025), Time-CMA (Liu et al. 2025), TQNet (Lin et al. 2025), Time-LLM (Jin et al. 2024), LLM-TS (Chen et al. 2025), iTransformer (Liu et al. 2024b), DLinear (Zeng et al. 2023), Time-VLM (Zhong et al. 2025), S2IP-LLM (Pan et al. 2024a), PatchTST (Nie et al. 2023), TimesNet (Wu et al. 2023), FEDFormer (Yang et al. 2024b), Informer (Zhou et al. 2021), Reformer (Kitaev, Kaiser, and Levskaya 2020), Koopa (Liu et al. 2023), and FPT (Zhou et al. 2023), some of which serve as supplementary references.
Comparative Results
Long Term Forecasting. The full results for long-term series forecasting are shown in Table 9. On the ETTh1, weather, Traffic and Solar datasets, we achieve the best performance. On the Weather datasets, we rank second MAE to TimeMixer++ and Time-LLM. Overall, our performance surpasses current SOTA methods.
Models SE-LLM TimeMixer++ FSCA Time-VLM AutoTimes S2IP-LLM Time-LLM FPT iTransformer
2025 2025 2025 2024 2024 2024 2024 2024
Yearly SMAPE 13.294 13.397 13.288 13.419 13.319 13.413 13.419 15.531 13.652
MASE 2.970 2.990 2.974 3.005 3.021 3.024 3.005 3.015 3.095
OWA 0.780 0.786 0.781 0.789 0.789 0.792 0.789 0.793 0.807
Quarterly SMAPE 10.079 10.206 10.037 10.110 10.020 10.352 10.110 10.177 10.353
MASE 1.177 1.201 1.174 1.178 1.162 1.228 1.178 1.194 1.209
OWA 0.887 0.901 0.884 0.889 0.878 0.922 0.889 0.897 0.911
Monthly SMAPE 12.618 12.720 12.762 12.980 12.696 12.995 12.980 12.894 13.079
MASE 0.931 0.943 0.947 0.963 0.936 0.970 0.963 0.956 0.974
OWA 0.875 0.884 0.897 0.903 0.880 0.910 0.903 0.897 0.911
Others SMAPE 4.896 4.593 4.761 4.795 4.916 4.805 4.795 4.940 4.780
MASE 3.306 3.380 3.207 3.178 3.310 3.247 3.178 3.228 3.231
OWA 1.037 1.054 1.007 1.006 1.039 1.017 1.006 1.029 1.012
Average SMAPE 11.778 11.884 11.828 11.894 11.831 12.021 11.983 11.991 12.142
MASE 1.578 1.597 1.580 1.592 1.585 1.612 1.595 1.600 1.631
OWA 0.847 0.856 0.850 0.855 0.850 0.857 0.859 0.861 0.874
Others 6.269 6.259 7.081 9.194 7.087 6.863 7.529 7.348 7.001
Avg 13.024 13.036 13.125 15.337 13.228 14.553 15.047 19.047 14.092
Quarterly 9.011 9.350 10.130 9.740 9.620 12.260 9.480 13.000 9.760
Others 4.872 5.790 4.810 6.810 9.440 6.880 8.940 13.030 7.530
Avg 12.560 12.750 13.060 14.030 13.390 14.170 13.530 15.820 13.370
Short Term Forecasting. The full results for short-term series forecasting are shown in Table 10. The experimental results indicate that We achieved the best performance in Yearly and Monthly patterns. In the Others pattern, Time-Mixer++ achieved the lowest SMAPE, while AutoTimes reached the lowest error in Quarterly pattern. From the lowest average error overall, SE-LLM achieve a lower SMAPE, signifying the best forecasting performance.
Zero-Shot Forecasting. The full results for zero-shot forecasting are shown in Table 11. We adopt the zero-shot forecasting methodology from AutoTimes, where each experimental setup involves training a model on a designated source dataset and subsequently deploying it to generate forecasts on a target dataset. These zero-shot scenarios are executed across subsets categorized by their sampling frequencies, each exhibiting distinct distributional properties. Overall, SE-LLM achieves the best performance.
Ablation Study
Ablation on Long-Term Forecasting. The full results of the innovation methods are shown in Table 12. SE-LLM uses cross-domain attention alignment (Jin et al. 2024; Chang et al. 2025; Liu et al. 2025) as the baseline. After replacing it with TSCC, errors were reduced across different LLMs. To make LLMs more suitable for processing time-series data, we incorporated a Time-Adapter, which further improved the model’s predictive performance. The ablation results demonstrate the effectiveness of our innovations.
LLM GPT2 Bert Opt125m Qwen2.5-0.5B
Dataset ECL Traffic ECL Traffic ECL Traffic ECL Traffic
Mertic MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
Baseline 96 0.145 0.240 0.387 0.267 0.150 0.245 0.440 0.304 0.140 0.237 0.368 0.258 0.135 0.233 0.373 0.263
192 0.165 0.257 0.410 0.277 0.168 0.262 0.459 0.314 0.158 0.254 0.390 0.268 0.152 0.249 0.394 0.272
336 0.183 0.275 0.427 0.286 0.187 0.281 0.476 0.322 0.176 0.272 0.407 0.292 0.170 0.267 0.409 0.280
720 0.224 0.309 0.463 0.306 0.230 0.316 0.514 0.344 0.216 0.305 0.445 0.300 0.212 0.303 0.444 0.300
Avg 0.179 0.270 0.422 0.284 0.184 0.276 0.472 0.321 0.173 0.267 0.403 0.280 0.167 0.263 0.405 0.279
+TSCC 96 0.138 0.233 0.368 0.257 0.144 0.240 0.395 0.277 0.136 0.232 0.361 0.251 0.134 0.231 0.356 0.247
192 0.157 0.250 0.393 0.269 0.161 0.256 0.417 0.286 0.154 0.249 0.383 0.263 0.152 0.248 0.377 0.257
336 0.176 0.269 0.414 0.281 0.181 0.274 0.436 0.295 0.173 0.267 0.401 0.271 0.169 0.266 0.393 0.266
720 0.218 0.305 0.461 0.312 0.223 0.309 0.474 0.316 0.215 0.303 0.439 0.294 0.210 0.301 0.431 0.287
Avg 0.172 0.264 0.409 0.280 0.177 0.270 0.431 0.294 0.170 0.263 0.396 0.270 0.166 0.262 0.389 0.264
+Time-Adapter 96 0.133 0.227 0.363 0.255 0.135 0.230 0.378 0.266 0.132 0.227 0.354 0.247 0.130 0.225 0.350 0.243
192 0.152 0.244 0.388 0.267 0.152 0.247 0.400 0.275 0.149 0.243 0.378 0.258 0.148 0.242 0.373 0.254
336 0.169 0.262 0.411 0.281 0.170 0.265 0.417 0.283 0.166 0.261 0.396 0.267 0.164 0.259 0.390 0.263
720 0.211 0.299 0.461 0.314 0.209 0.298 0.454 0.303 0.204 0.295 0.434 0.291 0.203 0.293 0.429 0.285
Avg 0.166 0.258 0.406 0.279 0.167 0.260 0.412 0.282 0.163 0.257 0.391 0.266 0.161 0.255 0.386 0.261
Ablation on Short-Term Forecasting. The ablation experiments for short-term forecasting are presented in the Table 13. We compared the performance by replacing different Large Language Models (LLMs) with SE-LLM on the M4 dataset. The findings indicate that LLama exhibits lower errors, followed by Qwen2.5-0.5B.
Table 13: The first column represents different LLMs replaced based on SE-LLM. ”Others” is obtained by averaging the results from the weekly, daily, and hourly periods. The values in bold represent the average results, while red indicates the best performance, followed by blue.
LLM GPT2 Bert Opt-125m Qwen2.5 Llama
sMAPE 13.843 15.659 14.028 13.396 13.294
Year MASE 3.199 3.674 3.198 3.002 2.970
OWA 0.826 0.941 0.832 0.788 0.780
sMAPE 10.300 12.237 10.320 10.168 10.079
Quarterly MASE 1.198 1.573 1.207 1.182 1.177
OWA 0.904 1.129 0.909 0.893 0.887
sMAPE 12.935 14.376 13.273 12.738 12.618
Monthly MASE 0.957 1.118 1.021 0.938 0.931
OWA 0.898 1.024 0.940 0.883 0.875
sMAPE 5.061 6.358 5.191 4.998 4.896
Others MASE 3.468 4.367 3.507 3.417 3.306
OWA 1.079 1.358 1.099 1.065 1.037
sMAPE 12.121 13.757 12.334 11.886 11.778
Average MASE 1.660 1.977 1.691 1.595 1.578
OWA 0.881 1.024 0.897 0.855 0.847
Table 14: Time-Adapter compared with LoRA. The input length is 672 and forecasting horizons were set to $\{96,192,336,720\}$.
@@
Metrics MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
TSCC 96 0.352 0.393 0.162 0.245 0.356 0.247 0.134 0.231 0.175 0.230
192 0.378 0.411 0.209 0.259 0.377 0.257 0.152 0.248 0.194 0.245
336 0.391 0.420 0.263 0.299 0.393 0.266 0.169 0.266 0.209 0.257
720 0.402 0.437 0.332 0.346 0.431 0.287 0.210 0.301 0.227 0.270
Avg 0.381 0.415 0.242 0.287 0.389 0.264 0.166 0.262 0.201 0.251
+LoRA 96 0.362 0.399 0.157 0.210 0.357 0.251 0.132 0.228 0.173 0.228
192 0.393 0.420 0.204 0.254 0.378 0.260 0.149 0.244 0.194 0.245
336 0.410 0.432 0.260 0.297 0.394 0.269 0.166 0.261 0.210 0.260
720 0.418 0.447 0.339 0.351 0.430 0.290 0.205 0.295 0.229 0.277
Avg 0.396 0.425 0.240 0.278 0.390 0.268 0.163 0.257 0.202 0.253
+Time-Adapter 96 0.355 0.398 0.150 0.200 0.350 0.243 0.130 0.225 0.165 0.220
192 0.389 0.421 0.194 0.243 0.373 0.254 0.148 0.242 0.183 0.236
336 0.409 0.435 0.247 0.285 0.390 0.263 0.164 0.259 0.199 0.249
720 0.418 0.448 0.323 0.339 0.429 0.285 0.203 0.293 0.219 0.266
Avg 0.393 0.426 0.229 0.267 0.386 0.261 0.161 0.255 0.192 0.243
Ablation on Adapters. We use the best-performing Qwen2.5-0.5B model as the baseline and evaluate the proposed Time-Adapter’s performance on time-series forecasting tasks. The results are compared to the LoRA adapter. The LoRA parameter configuration is kept consistent with the Time-Adapter, with a rank of 8 and the full ablation results are presented in Table 14. The comparison protocol involves evaluating both adapters across several time-series datasets, specifically focusing on the performance metrics for each dataset. Our findings reveal that the LoRA adapter shows improvements in only the weather and ECL datasets, whereas the Time-Adapter consistently enhances performance across all datasets. In particular, the error reduction is most notable in the weather dataset, where the Time-Adapter achieves a 5.4% decrease in MSE compared to the baseline. This suggests that the Time-Adapter provides a more generalized performance boost across diverse time-series data, as compared to LoRA.